OpenStudy (anonymous):

the sum of two consecutive integers is -211 find the integers

OpenStudy (gudden):

let one interger be x then the other = x +1 ( since the intergers are consecutive) x + x + 1 =-211 2x +1 = - 211 on adding -1 from both sides we get 2x = -212 and on dividing the equation wth 2 we get x = - 106 and the other = - 105

OpenStudy (anonymous):

Let's call one integer a and the other b. b will be the larger of the two, so b=a+1. You know that a+b=-211. Substitute a+1 in for b... a+a+1=-211, which can be simplified: 2a+1=-211, and then it's basic algebra. 2a+1-1=-211-1 2a=-212 a=-106 b=a+1 b=-106+1 b=-105 Let's verify that. Does a+b=-211? a+b=-211 -106+-105=-211 -211=-211 So your answer is -106 and -105.

OpenStudy (anonymous):

-211/2 =-105.5 -105 and -106